A bus station has exactly four entrances. In the morning the numbers of passengers arriving at these entrances during a 10-second period have the independent distributions Po(0.4), Po(0.1), Po(0.2) and Po(0.5). Find the probability that the total number of passengers arriving at the four entrances to the bus station during a randomly chosen 1-minute period in the morning is more than 3.
- Find the probability that the total number of passengers arriving at the four entrances to the bus station during a randomly chosen 1-minute period in the morning is more than 3.[3]
Mark scheme
Marking points from the official mark scheme (May/June 2024). The mark or mark type is shown in brackets.
- [B1] λ = 7.2
- [M1] P(X> 3) = 1 - e-7.2(1+7.2+7.2^2/2!+7.2^3/3!) or 1 - e-7.2(1 + 7.2 + 25.92 + 62.21) or 1 - (0.0007466 + 0.005375 +0.01935 + 0.04644)
- [A1] = 0.928 (3sf)
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →